Find three consecutive even integers, such that 7 times the middle integer is 10 more than the sum of first and third.
Answer by mananth(16946)

let x,x+2,x+4 be the integers
7(x+2)= x+x+4+10
7x+14=2x+14
5x=0
x=0
0,2,4